# Dissecting the SWE-Bench Leaderboards: Profiling Submitters and Architectures of LLM- and Agent-Based Repair Systems

### v1

@@ 3.4.2. Issue Reproduction
Submissions such as SWE-Agent and CodeR leverage LLMs to generate reproduction tests that aim to replicate the original issue directly from issue descriptions. Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) generates multiple candidate reproduction tests and selects the optimal one based on actual execution results on the original codebase. SWE-RL (wei2025SWE-RL, ) builds upon this approach through the creation of the agent scaffold Agentless Mini, introducing two key improvements: 1) Instead of relying solely on the issue description, it retrieves a relevant existing test file to guide test generation; and 2) Rather than selecting a single majority-voted test, it allows the selection of multiple top-ranked test samples based on voting outcomes. PatchPilot extends Agentless with a self-reflection-based PoC reproduction mechanism, inspired by the Reflexion mechanism (shinn2023reflexion, ). To address common issues at this stage—such as misconfigurations and missing dependencies—the approach leverages an LLM to iteratively generate and refine scripts that reproduce the issue (referred to as a Proof-of-Concept, or PoC). It uses carefully crafted prompts to ensure the inclusion of necessary dependencies and configurations.
@@
Agents specialized in issue reproduction are incorporated in various multi-agent approaches. For instance, the Bytedance MarsCode Agent features a Reproducer Agent that generates reproduction scripts based on the relevant code and issue description, and performs dynamic debugging in a sandbox to confirm successful reproduction (liu2024marscode, ). SpecRover also includes a Reproducer Agent that writes a test to reproduce the program fault described in the issue, which is also used to verify the patch. Since the test is generated by an LLM and may vary due to its non-deterministic nature, SpecRover further employs a Reviewer Agent to assess the correctness of the test based on the issue description.
@@
One of the primary obstacles in issue reproduction is the creation of a runnable environment, as emphasized by SWE-RL (wei2025SWE-RL, ). To address this, the multi-agent system MASAI (arora2024masaimodulararchitecturesSE, ) defines two specialized sub-agents for this task: 1) the Test Template Generator, which analyzes documentation, existing tests, and repository setup to determine how to write and execute new tests. It then produces a repository-specific, issue-independent test template along with the command required to run it; and 2) the Issue Reproducer, which uses the generated test template to write a test that replicates the behavior described in the issue. nFactorial defines a Reproducer Agent tasked with generating a standalone Python script to replicate the issue, using tools for executing bash scripts and editing files, and ensuring the result is reliably reproducible across multiple executions. DARS (aggarwal2025darsdynamicactionresampling, ) equips agents with actions for creating and improving reproduction scripts based on previous attempts.
@@
Notably, some approaches have incorporated issue reproduction progressively over time. For example, GRU has two submissions on each leaderboard, and one key difference between them is the addition of a new task in the workflow aimed at a reproduction script.
@@ 3.4.3. Issue Localization
The goal of issue localization is to identify code elements —at a specific granularity such as line— that are suspected to be related to the issue and may require modification. We adopt the four-category taxonomy proposed by Liu et al. (liu2024SurveyAgentsLLM4SE, ), which classifies seven agent-based approaches according to their localization strategies.
@@
The use of retrieval-based strategies enables approaches to identify relevant code elements by measuring their similarity to inputs such as issue descriptions. Our analysis reveals that multiple submissions rely on large language models (LLMs) to perform at least some steps of the localization process. For example, SWE-Fixer (xie2025SWEFixerTrainingopensourcellms, ) adopts a coarse-to-fine strategy, first applying BM25 (RobertsonBM25, ) for initial file retrieval, and then using a model to identify the defective files from the retrieved set.
Repository representations created during the preprocessing phase (Section 3.4.1) are leveraged in the localization phase alongside LLMs. For instance, DARS employs a keyword-based subgraph retrieval technique (ego-graphs (hu2024graggraphretrievalaugmentedgeneration, )) over its RepoGraph, enabling it to identify files and specific line numbers relevant to the issue. Composio SWE-Kit includes a tool that analyzes the codebase and create a structured code map. This map allows the agent to efficiently locate, retrieve, and understand relevant code sections without -according to the authors- overloading the LLM’s context. A similar strategy, combining a repository map with keyword-based search, is employed by CodeStory. Amazon Q Developer Agent leverages a structured XML representation of the repository, along with additional context, to guide an LLM in identifying relevant files for retrieval. SuperCoderAGI employs a hierarchical search space reduction strategy, using an LLM at each step to progressively narrow down from candidate files to specific locations within those files, via RAG and repository maps. SWE-RL (wei2025SWE-RL, ), through its scaffold Agentless Mini, uses an LLM to predict relevant file paths based on the given issue and a repository structure mapped to file paths. HyperAgent incorporates an LLM-based agent named Navigator, specialized in information retrieval within the repositiry.
In contrast, some approaches do not rely on LLMs for localization. For example, AppMap Navie defines a method to retrieve the relevant code context based on the issue text. According to the authors, this method operates locally without invoking an LLM, making it fast, predictable, and efficient.
@@
The navigation strategy allows agents to perform code search actions across the entire repository structure (liu2024SurveyAgentsLLM4SE, ). Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) employs a hierarchical localization process that first narrows down the fault to specific files, then to relevant classes or functions, and finally to fine-grained edit locations. This is achieved through a combination of LLM-based techniques and classic information retrieval-based localization methods (Zhou2012RetrievalBug, ). Both SWE-Agent (yang2024sweagentagent, ) and MASAI (arora2024masaimodulararchitecturesSE, ) include tools that enable their agents to navigate the repository to locate relevant files or to visualize files to extract code elements. AutoCodeRover (Zhang2024AutoCoderRover, ) uses an LLM-based agent, Context Retrieval, to navigate the codebase and extract relevant snippets, leveraging context retrieval APIs such as search_method_in_class which works at the AST-level.
@@
Agents dedicated to localization are also employed. For instance, IBM AI Agent SWE-1.0 employs multiple localization agents, each potentially backed by a different LLM, that collaborate with a judge to identify the bug location (file, function, and line) by following a Think-Act-Observe loop and utilizing tools based on BM25, LLM-based search and syntactic search. nFactorial includes a Localizer agent designed to navigate a custom file tree structure and identify code segments relevant to the problem statement. Similarly, AugmentCode equips its agents with tools such as grep and find for repository navigation. Globant Code Fixer Agent features a Localization stage, where a set of agents explores the codebase to identify candidate buggy.
@@
The structured representation of a repository (Section 3.4.1) is used to support navigation. For example, Agentless + RepoGraph extends Agentless by integrating RepoGraph, a structured repository graph that supports keyword-based queries (e.g., issue derived) through the search_repograph() method. which searches on the build graph given as input key terms, such as some extracted from the issue description. Nemotron-Cortexa employs a Localization Agent in a two-step process: first, it identifies relevant files using a custom-trained code embedding model aligned with issue descriptions; then, through candidate entity filtering, the agent iteratively traverses the code graph to refine results at the level of functions, classes, or methods. Aider uses the built repository map, compact and powerful summary of the entire codebase, to help the LLM to decide which file to edit, and a justification of that decision. This map is constantly tailored show repository context that is relevant to the current state of the problem. Aider uses a repository map —a compact summary of the entire codebase— to help the LLM decide which file to edit and provide a justification, continuously tailoring the map to reflect the context relevant to the current problem state. ORCALOCA (yu2025orcaLocallmAgent, ) enhances issue localization with LLM-based agents by combining dynamic action scheduling, hierarchical action decomposition, and a context manager that prunes irrelevant information to ensure focused and efficient codebase exploration.
@@
Spectrum-based fault localization (SBFL) techniques have been widely used in traditional APR systems, such as GenProg (LeGoues2011genprog, ), to compute the suspiciousness of code elements (e.g., lines) based on their coverage in failing and passing test cases. AutoCodeRover uses SBFL to augment the problem statement (i.e., the issue description) with additional relevant classes and methods. CodeR adopts a hybrid approach that linearly combines suspiciousness scores from SBFL (based on tests generated by its Reproducer agents; see Section 3.4.2) and BM25-based file-level localization using issue similarity. Unlike SBFL-based methods, PatchPilot (li2025patchpilot, ) adopts a runtime-driven hybrid approach that executes the issue to collect execution traces, uses LLMs to filter relevant files and perform fine-grained searches, and finally applies an LLM-based review step to retrieve code snippets linked to the root cause.
@@
RepoUnderstander (OLDma2024understandRepoUnderstander, ), later renamed Alibaba Lingma Agent(ma2024lingmaswegpt, ), performs issue localization through simulation. It first constructs a Repository Knowledge Graph (see Section LABEL:) and then applies a Monte Carlo Tree Search (MCTS) strategy to explore the graph. By simulating multiple exploration trajectories and assigning reward values to graph nodes (i.e., code elements), the approach narrows the search space to localize faults more precisely. MCTS has also been explored by other approaches, such as CodeStory Midwit Agent. However, according to the submission authors, they dropped MCTS from their framework due to long execution times.
@@ 3.4.4. Task Decomposition - Planning
There are approaches that decompose the the main task (in this case, resolve an issue) into more fine-grain sub-tasks. LingmaAgent (ma2025alibabalingmaagent, ) includes an LLM-based agent, Summary Agent, which takes as input the relevant code fragments identified through MCTS-based repository understanding along with the issue description, and produces a sequential summary of the relevant code and plans a solution. AppMap Navie performs a planning step by combining the issue description with context to produce a structured plan outlining the problem, a high-level solution, target files, and file-specific change descriptions without code. Similarly, Gru takes the list of relevant files from localization, analyzes the corresponding code, and decides which files to modify and how the modifications should be made.
The submissions Tools + Claude by Anthropic equip an LLM (Claude 3.5/3.7 Sonnet) with a set of tools. In addition to the Edit and Bash tools, it introduces a “planning” tool that prompts the LLM to write down its thoughts as it solves the problem. This tool leverages the inherent reasoning capabilities of Claude Sonnet’s default thinking mode. As the details of the planning tool were not publicly disclosed, the submission Augment Agent replaces it with the Sequential Thinking MCP framework, which enables dynamic and reflective problem-solving through a structured reasoning process.
Agents for plan elaboration have been employed. For example, CodeR (xia2024dcoder, ) is a multi-agent system featuring a Manager Agent responsible for selecting one of four predefined plans based on the issue description and interpreting the execution summary after the plan is carried out. A similar approach is adopted by MarsCode (liu2024marscode, ), which includes a Planner that analyzes the collected code snippets and classifies the issue into one of two workflows: dynamic debugging repair or static repair. SuperCoder (gautam2024supercoder, ) includes a PlannerAgent responsible for navigating code files, identifying buggy locations, and determining the changes needed to resolve the issues. HyperAgent (phan2024hyperagent, ) is a multi-agent system centered around a Planner agent that processes issues, generates resolution strategies, and coordinates specialized child agents (Navigator, Editor, Executor) via an asynchronous message queue in an iterative planning and feedback loop.
Other submissions, such as Aegis, Amazon Q Developer Agent, and Google Jules, mention the inclusion of a planning phase or the generation of a plan, though they do not provide further implementation details.

### v2

@@ 3.3.2. Issue Reproduction
Submissions such as SWE-Agent (56.67% on Lite, 66.6% on Verified) and CodeR (28.33% on Lite) leverage LLMs to generate reproduction tests that aim to replicate the original issue directly from issue descriptions. Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) (40.67% on Lite, 50.8% on Verified) generates multiple candidate reproduction tests and selects the optimal one based on actual execution results on the original codebase. SWE-RL (wei2025SWE-RL, ) (41.2% on Verified) builds upon this approach through the creation of the Agentless Mini agent scaffold, introducing two key improvements: 1) Instead of relying solely on the issue description, it retrieves a relevant existing test file to guide test generation; and 2) Rather than selecting a single majority-voted test, it allows the selection of multiple top-ranked test samples based on voting outcomes. PatchPilot (41.33% on Lite, 64.6% on Verified) extends Agentless with a self-reflection-based PoC reproduction mechanism, inspired by the Reflexion mechanism (shinn2023reflexion, ). To address common issues at this stage —such as misconfigurations and missing dependencies— the approach leverages an LLM to iteratively generate and refine scripts that reproduce the issue (referred to as a Proof-of-Concept, or PoC). It uses carefully crafted prompts to ensure the inclusion of necessary dependencies and configurations.
@@
Agents specialized in issue reproduction are incorporated in various multi-agent approaches. For instance, Bytedance MarsCode Agent (39.33% on Lite, 50% on Verified) features a Reproducer Agent that generates reproduction scripts based on the relevant code and issue description, and performs dynamic debugging in a sandbox to confirm successful reproduction (liu2024marscode, ). SpecRover (51.6% on Verified) also includes a Reproducer Agent that writes a test to reproduce the program fault described in the issue, which is also used to verify the patch. Since the test is generated by an LLM and may vary due to its non-deterministic nature, SpecRover further employs a Reviewer Agent to assess the correctness of the test based on the issue description.
@@
One of the primary obstacles in issue reproduction is the creation of a runnable environment, as emphasized by SWE-RL (wei2025SWE-RL, ) (41.2% on Verified). To address this, the multi-agent system MASAI (arora2024masaimodulararchitecturesSE, ) (27.33% on Lite, 32.6% on Verified) defines two specialized sub-agents for this task: 1) the Test Template Generator, which analyzes documentation, existing tests, and repository setup to determine how to write and execute new tests. It then produces a repository-specific, issue-independent test template along with the command required to run it; and the Issue Reproducer, which uses the generated test template to write a test that replicates the behavior described in the issue. nFactorial (49.2% on Verified) defines a Reproducer Agent tasked with generating a standalone Python script to replicate the issue, using tools for executing bash scripts and editing files, and ensuring the result is reliably reproducible across multiple executions. DARS (aggarwal2025darsdynamicactionresampling, ) (47% on Lite) equips agents with actions for creating and improving reproduction scripts based on previous attempts.
@@
Some approaches such as Alibaba Lingma Agent (33% on Lite, 28.8% on Verified) and SWE-fixer (24.67% on Lite, 32.8% on Verified) do not offer issue reproduction. Notably, some others have incorporated issue reproduction progressively over time. For example, GRU has two submissions on each leaderboard, and one key difference between them is the addition of a new task in the workflow aimed at a reproduction script.
@@ 3.3.3. Issue Localization
The goal of issue localization is to identify code elements —at a specific granularity, such as line— that are suspected to be related to the issue and may require modification. We adopt the four-category taxonomy proposed by Liu et al. (liu2024SurveyAgentsLLM4SE, ), which classifies seven agent-based approaches according to their localization strategies.
@@
The use of retrieval-based strategies enables approaches to identify relevant code elements by measuring their similarity to inputs such as issue descriptions. Several techniques are used:
Our analysis reveals that multiple submissions rely on large language models (LLMs) to perform at least some steps of the localization process. For example, SWE-Fixer (xie2025SWEFixerTrainingopensourcellms, ) (24.67% on Lite, 32.8% on Verified) adopts a coarse-to-fine strategy, first applying BM25 (RobertsonBM25, ) for initial file retrieval, and then using a model to identify the defective files from the retrieved set.
Repository representations created during the preprocessing phase (Section 3.3.1) are leveraged in the localization phase alongside LLMs. For instance, DARS (47% on Lite) employs a keyword-based subgraph retrieval technique (ego-graphs (hu2024graggraphretrievalaugmentedgeneration, )) over its RepoGraph, enabling it to identify files and specific line numbers relevant to the issue. Composio SWE-Kit (41% on Lite, 48.6% on Verified) includes a tool that analyzes the codebase and create a structured code map. This map allows the agent to efficiently locate, retrieve, and understand relevant code sections without -according to the authors- overloading the LLM’s context. A similar strategy, combining a repository map with keyword-based search, is employed by CodeStory.
KGCompass (yang2025enhancing, ) (46% on Lite) employs a hybrid approach that combines knowledge graph analysis with LLM-based reasoning to identify potential bug locations. It first ranks function entities based on their relevance to the issue description, using a score that integrates semantic and textual similarity to the issue text as well as structural proximity in the graph; then it selects the top 15 candidates. An LLM further analyzes the problem description to identify up to 5 additional locations.
Amazon Q Developer Agent (29.67% on Lite, 65.4% on Verified) leverages a structured XML representation of the repository, along with additional context, to guide an LLM in identifying relevant files for retrieval. SuperCoder  (34% on Lite) employs a hierarchical search space reduction strategy, using an LLM at each step to progressively narrow down from candidate files to specific locations within those files, via RAG and repository maps. SWE-RL (wei2025SWE-RL, ) (41.2% on Verified), through its scaffold Agentless Mini, uses an LLM to predict relevant file paths based on the given issue and a repository structure mapped to file paths. HyperAgent (25.33% on Lite) incorporates an LLM-based agent named Navigator, specialized in information retrieval within the repository.
In contrast, some approaches do not rely on LLMs for localization. For example, AppMap Navie (36% on Lite, 47.2% on Verified) defines a method to retrieve the relevant code context based on the issue text. According to the authors, this method operates locally without invoking an LLM, making it fast, predictable, and efficient.
@@
The navigation strategy allows agents to perform code search actions across the entire repository structure (liu2024SurveyAgentsLLM4SE, ). Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) (40.67% on Lite, 50.8% on Verified) employs a hierarchical localization process that first narrows down the fault to specific files, then to relevant classes or functions, and finally to fine-grained edit locations. This is achieved through a combination of LLM-based techniques and classic information retrieval-based localization methods (Zhou2012RetrievalBug, ). Both SWE-Agent (yang2024sweagentagent, ) (56.67% on Lite, 66.6% on Verified) and MASAI (arora2024masaimodulararchitecturesSE, ) (27.33% on Lite, 32.6% on Verified) include tools that enable their agents to navigate the repository to locate relevant files or to visualize files to extract code elements.
Co-PatcherR (tang2025copatchercollaborativesoftwarepatching, ) (46% on Verified), build on top of PatchPilot (li2025patchpilot, ), improves localization by dividing it into two small-LLM-driven sub-tasks: file localization, where an LLM receives the issue description and repository file structure to identify relevant files, and line localization, where the LLM analyzes code from each selected file along with the issue description to pinpoint faulty lines. Using Qwen-2.5-Coder as the base model, the localization LLM is fine-tuned on distillation data with reasoning chains generated by Claude-3.7-Sonnet.
@@
For instance, IBM AI Agent SWE-1.0 (23.67% on Lite) employs multiple localization agents, each potentially backed by a different LLM, that collaborate with a judge to identify the bug location (file, function, and line) by following a Think-Act-Observe loop and utilizing tools based on BM25, LLM-based search and syntactic search. AutoCodeRover (Zhang2024AutoCoderRover, ) (38.4% on Lite, 30.67% on Verified) uses an LLM-based agent, Context Retrieval, to navigate the codebase and extract relevant snippets, leveraging context retrieval APIs such as search_method_in_class which works at the AST-level. Nemotron-CORTEXA (sohrabizadeh2025nemotroncortexa, ) (68.2% on Verified) employs a Localization Agent in a two-step process: first, it identifies relevant files using a custom-trained code embedding model aligned with issue descriptions; then, through candidate entity filtering, the agent iteratively traverses the code graph to refine results at the level of functions, classes, or methods. nFactorial (49.2% on Verified) includes a Localizer agent designed to navigate a custom file tree structure and identify code segments relevant to the problem statement. Similarly, Augment Agent (70.4% on Verified) equips its agents with tools such as grep and find for repository navigation. Globant Code Fixer Agent (48.33% on Lite) features a Localization stage, where a set of agents explores the codebase to identify candidate buggy. ORCALOCA (yu2025orcaLocallmAgent, ) (41% on Lite) enhances issue localization with LLM-based agents by combining dynamic action scheduling, hierarchical action decomposition, and a context manager that prunes irrelevant information to ensure focused and efficient codebase exploration.
@@
The structured representation of a repository (Section 3.3.1) is used to support navigation. For example, Agentless + RepoGraph (29.67% on Lite) extends Agentless by integrating RepoGraph, a structured repository graph that supports keyword-based queries (e.g., issue derived) through the search_repograph() method. which searches on the build graph given as input key terms, such as some extracted from the issue description. Aider (26.33% on Lite) uses the built repository map, compact and powerful summary of the entire codebase, to help the LLM to decide which file to edit, and a justification of that decision. This map is constantly tailored show repository context that is relevant to the current state of the problem.
@@
Spectrum-based fault localization (SBFL) techniques have been widely used in traditional APR systems, such as GenProg (LeGoues2011genprog, ), to compute the suspiciousness of code elements (e.g., lines) based on their coverage in failing and passing test cases. AutoCodeRover (38.4% on Lite, 30.67% on Verified) complements the technique mentioned above with SBFL in order to augment the problem statement (i.e., the issue description) with additional relevant classes and methods. CodeR (28.33% on Lite) adopts a hybrid approach that linearly combines suspiciousness scores from SBFL (based on tests generated by its Reproducer agents; see Section 3.3.2) and BM25-based file-level localization using issue similarity. Unlike SBFL-based methods, PatchPilot (li2025patchpilot, ) adopts a runtime-driven hybrid approach that executes the issue to collect execution traces, uses LLMs to filter relevant files and perform fine-grained searches, and finally applies an LLM-based review step to retrieve code snippets linked to the root cause.
@@
Alibaba Lingma Agent(ma2024lingmaswegpt, ) (33% on Lite, 28.8% on Verified) performs issue localization through simulation. It first constructs a Repository Knowledge Graph (see Section 3.3.1) and then applies a Monte Carlo Tree Search (MCTS) strategy to explore the graph. By simulating multiple exploration trajectories and assigning reward values to graph nodes (i.e., code elements), the approach narrows the search space to localize faults more precisely. MCTS has also been explored by other approaches, such as CodeStory Midwit Agent (62.2% on Verified). However, according to the submission authors, they dropped MCTS from their framework due to long execution times.
Combining techniques
Codev identifies relevant files using a multi-source approach that combines static analysis, semantic search, and LLM-based reasoning. Six complementary sources are used to retrieve potentially relevant code blocks: a code graph, two types of string search, a semantic embedding model, an LLM-based analyst, and an agentic top-down search. Each retrieved block is evaluated and scored by a lightweight LLM, which classifies them as Solution Target, Supporting, Interesting, or Not Relevant. These scores are aggregated into a confidence measure per file, guiding the selection of files for subsequent development planning.
@@ 3.3.4. Task Decomposition - Planning
There are approaches that decompose the main task (in this case, resolve an issue) into more fine-grain sub-tasks. The Tools + Claude entries by Anthropic (73.2% on Verified) furnish an LLM (Claude 3.5/3.7/4 Sonnet) with a set of tools for planning. In addition to the Edit and Bash tools, they introduces a “planning” tool that prompts the LLM to write down its thoughts as it solves the problem. This tool leverages the inherent reasoning capabilities of Claude Sonnet’s default thinking mode. As the details of the planning tool were not publicly disclosed, the submission Augment Agent (70.4% on Verified) replaces it with the Sequential Thinking MCP framework, which enables dynamic and reflective problem-solving through a structured reasoning process.
Agents for plan elaboration are also employed. For example, Alibaba Lingma Agent (ma2025alibabalingmaagent, ) (33% on Lite, 28.8% on Verified) includes an LLM-based agent, Summary Agent, which takes as input the relevant code fragments identified through MCTS-based repository understanding along with the issue description, and produces a sequential summary of the relevant code and plans a solution. CodeR (xia2024dcoder, ) (28.33% on Lite) is a multi-agent system featuring a Manager Agent responsible for selecting one of four predefined plans based on the issue description and interpreting the execution summary after the plan is carried out. A similar approach is adopted by Bytedance MarsCode (liu2024marscode, ) (39.33% on Lite, 50% on Verified), which includes a Planner that analyzes the collected code snippets and classifies the issue into one of two workflows: dynamic debugging repair or static repair. SuperCoder (gautam2024supercoder, ) (34% on Lite) includes a PlannerAgent responsible for navigating code files, identifying buggy locations, and determining the changes needed to resolve the issues. HyperAgent (phan2024hyperagent, ) (25.33% on Lite) is a multi-agent system centered around a Planner agent that processes issues, generates resolution strategies, and coordinates specialized child agents (Navigator, Editor, Executor) via an asynchronous message queue in an iterative planning and feedback loop.
Some approaches do not detail the use of LLMs or agents. AppMap Navie performs a planning step by combining the issue description with context to produce a structured plan outlining the problem, a high-level solution, target files, and file-specific change descriptions without code. Similarly, Gru (48.67% on Lite, 57% on Verified) takes the list of relevant files from localization, analyzes the corresponding code, and decides which files to modify and how the modifications should be made. In the extreme case, approaches such as Aegis (30.33% on Lite), Amazon Q Developer Agent (29.67% on Lite, 65.4% on Verified), and Google Jules (52.2% on Verified), mention the inclusion of a planning phase or the generation of a plan, but they do not provide further implementation details.